1. A 550 g mass is tied to a string and made to move in a vertical circle of radius 55 cm at a constant speed. The tension in the string ranges from 2.0 N to 12.7 N. 

a. At what point is the tension a maximum? When is it a minimum?

b. Determine the speed of the mass.



1
Expert's answer
2021-11-11T12:02:44-0500

The tension is maximum in the lowest point.

Find the speed of the object.


"T_d=mg+mv^2/R,\\\space\\
v=\sqrt{R\bigg(\frac{T_d}{m}-g\bigg)}=1.7\text{ m/s}."